# Partial Rewriting and Value Interpretation of Logically Constrained Terms

### v1

@@ 4 Interpretations of Partial and Most General Constrained Reductions
    Let $u=s\hat{\delta}$ and let us show that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for some $v$. For this, it suffices to show $\ell\gamma\hat{\delta}\to^{\epsilon}_{\langle\!\langle\rho\rangle\!\rangle}r\gamma\hat{\delta}$ as $u|_{p}=(s\hat{\delta})|_{p}=s|_{p}\hat{\delta}=\ell\gamma\hat{\delta}$. Let $\sigma=(\hat{\delta}\circ\gamma)|_{Z\cap\mathcal{V}\mathsf{ar}(\ell,r)}$. It remains to show that $\mathcal{D}\mathsf{om}(\sigma)=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$ and $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
@@
    As $\mathcal{V}\mathsf{ar}(t)\cap\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}=\varnothing$, one can extend $\delta$ to $\delta^{\prime}=\delta\cup\{\vec{y}\mapsto\vec{v}\}$ so that $\vDash_{\delta^{\prime}}\psi$ for some $\vec{v}\in\mathcal{V}\mathsf{al}^{*}$. We further extend $\delta^{\prime}$ to $\delta^{\prime\prime}=\delta^{\prime}\cup\{\vec{x}\mkern 2.0mu\vphantom{x}\mapsto\vec{w}\}$ where $\{\hskip 0.85358pt\vec{x}\mkern 2.0mu\vphantom{x}\hskip 0.85358pt\}=X\setminus\mathcal{V}\mathsf{ar}(t)$ by taking an arbitrary $\vec{w}\in\mathcal{V}\mathsf{al}^{*}$.
    We now show $s\delta^{\prime\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. Firstly, by $Y=(X\cap\mathcal{V}\mathsf{ar}(t))\cup(\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell))$ and (10), we have $X\cap\mathcal{V}\mathsf{ar}(t)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta)$. Thus, $\delta^{\prime\prime}(X)\subseteq\mathcal{V}\mathsf{al}$, by our choice of $\delta^{\prime\prime}$. It remains to show $\delta^{\prime\prime}\vDash\exists\vec{x}.\ \varphi$. We have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{y}.\ \psi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta)$ by (11). Hence, $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq\mathcal{V}\mathsf{ar}(\varphi)\subseteq\mathcal{V}\mathsf{ar}(\psi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta^{\prime})\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta^{\prime\prime})$. Also, by $\vDash_{\delta^{\prime}}\psi\penalty 10000\ (=\varphi\land\pi\gamma)$, we have $\vDash_{\delta^{\prime}}\exists\vec{x}.\ \varphi$. Thus, since $\delta^{\prime}(x)=\delta^{\prime\prime}(x)$ for $x\notin\{\vec{x}\}$, we have $\vDash_{\delta^{\prime\prime}}\exists\vec{x}.\ \varphi$. Hence $\delta^{\prime\prime}\vDash\exists\vec{x}.\ \varphi$ as claimed. Thus, $s\delta^{\prime\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$.
@@
    As $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$, we have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$. Furthermore, from $\vDash_{\delta^{\prime}}\psi\penalty 10000\ (=\varphi\land\pi\gamma)$ and $\mathcal{V}\mathsf{ar}(\pi\gamma)\cap\{\vec{x}\}=\varnothing$, it follows that $\vDash_{\delta^{\prime\prime}}\pi\gamma$. Hence, we have $\vDash_{[\![\gamma]\!]_{\delta^{\prime\prime}}}\pi$. By $\mathcal{D}\mathsf{om}(\gamma)\cap\mathcal{V}\mathsf{ar}(\pi)\subseteq\mathcal{V}\mathsf{ar}(\ell)\cap X$, we have $\gamma]\!]_{\delta^{\prime\prime}}(x)=\delta^{\prime\prime}(\gamma(x))=\sigma(x)$ for all $x\in\mathcal{V}\mathsf{ar}(\pi)$. Hence, $\vDash_{\sigma}\pi$. Thus, $\vDash_{\sigma}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$, and $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ follows by $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$. This concludes the claim.
@@
Lemma˜4.8 immediately yields the following characterization of partial constrained reductions in terms of interpretation.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\stackrel{{\scriptstyle*}}{{\leadsto}}_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
  1. 1.
  2. 2.
Proof 4.11.
1 follows from item 2 of Lemma˜4.8. 2 follows from the items 1 and 3 of Lemma˜4.8.
Note that by Theorem˜3.6, the properties that hold for partial rewrite steps should hold also for most general rewrite steps. Thus, the properties presented in Theorem˜4.10 also hold for most general rewrite steps.
Lemma 4.12.
Proof 4.13.
@@
  • Otherwise, let $z\in\mathcal{V}\mathsf{ar}(\pi)\cap(\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell))$. Then, $\gamma]\!]_{\hat{\delta}^{\prime}}(z)=[\![\gamma]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}(z)=[\![\gamma(z)]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}=[\![z]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}=\{\vec{w}\mapsto\vec{v}\}(z)=\hat{\sigma}(z)$.

### v2

@@ Appendix A Omitted Proofs
    Let $u=s\hat{\delta}$ and let us show that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for some $v$. For this, it suffices to show $\ell\gamma\hat{\delta}\to^{\epsilon}_{\langle\!\langle\rho\rangle\!\rangle}r\gamma\hat{\delta}$ as $u|_{p}=(s\hat{\delta})|_{p}=s|_{p}\hat{\delta}=\ell\gamma\hat{\delta}$. Let $\sigma=(\hat{\delta}\circ\gamma)|_{Z\cap\mathcal{V}\mathsf{ar}(\ell,r)}$. It remains to show that $\mathcal{D}\mathsf{om}(\sigma)=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$ and $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
@@
    As $\mathcal{V}\mathsf{ar}(t)\cap\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}=\varnothing$, one can extend $\delta$ to $\delta^{\prime}=\delta\cup\{\hskip 0.85358pt\vec{y}\mapsto\vec{v}\hskip 0.85358pt\}$ so that $\vDash_{\delta^{\prime}}\psi$ for some $\vec{v}\in\mathcal{V}\mathsf{al}^{*}$. We further extend $\delta^{\prime}$ to $\delta^{\prime\prime}=\delta^{\prime}\cup\{\hskip 0.85358pt\vec{x}\mkern 2.0mu\vphantom{x}\mapsto\vec{w}\hskip 0.85358pt\}$ where $\{\hskip 0.85358pt\vec{x}\mkern 2.0mu\vphantom{x}\hskip 0.85358pt\}=X\setminus\mathcal{V}\mathsf{ar}(t)$ by taking an arbitrary $\vec{w}\in\mathcal{V}\mathsf{al}^{*}$.
    We now show $s\delta^{\prime\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$. Firstly, by $Y=(X\cap\mathcal{V}\mathsf{ar}(t))\cup(\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell))$ and (10), we have $X\cap\mathcal{V}\mathsf{ar}(t)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta)$. Thus, $\delta^{\prime\prime}(X)\subseteq\mathcal{V}\mathsf{al}$, by our choice of $\delta^{\prime\prime}$. It remains to show $\delta^{\prime\prime}\vDash\exists\vec{x}.\ \varphi$. We have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{y}.\ \psi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta)$ by (11). Hence, $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq\mathcal{V}\mathsf{ar}(\varphi)\subseteq\mathcal{V}\mathsf{ar}(\psi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta^{\prime})\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\delta^{\prime\prime})$. Also, by $\vDash_{\delta^{\prime}}\psi\penalty 10000\ (=\varphi\land\pi\gamma)$, we have $\vDash_{\delta^{\prime}}\exists\vec{x}.\ \varphi$. Thus, since $\delta^{\prime}(x)=\delta^{\prime\prime}(x)$ for $x\notin\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$, we have $\vDash_{\delta^{\prime\prime}}\exists\vec{x}.\ \varphi$. Hence $\delta^{\prime\prime}\vDash\exists\vec{x}.\ \varphi$ as claimed. Thus, $s\delta^{\prime\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$.
@@
    As $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq Z\cap\mathcal{V}\mathsf{ar}(\ell,r)$, we have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$. Furthermore, from $\vDash_{\delta^{\prime}}\psi\penalty 10000\ (=\varphi\land\pi\gamma)$ and $\mathcal{V}\mathsf{ar}(\pi\gamma)\cap\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}=\varnothing$, it follows that $\vDash_{\delta^{\prime\prime}}\pi\gamma$. Hence, we have $\vDash_{[\![\gamma]\!]_{\delta^{\prime\prime}}}\pi$. By $\mathcal{D}\mathsf{om}(\gamma)\cap\mathcal{V}\mathsf{ar}(\pi)\subseteq\mathcal{V}\mathsf{ar}(\ell)\cap X$, we have $\gamma]\!]_{\delta^{\prime\prime}}(x)=\delta^{\prime\prime}(\gamma(x))=\sigma(x)$ for all $x\in\mathcal{V}\mathsf{ar}(\pi)$. Hence, $\vDash_{\sigma}\pi$. Thus, $\vDash_{\sigma}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$, and $\sigma\vDash\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ follows by $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)$. This concludes the claim.
@@
See 4.8
Proof A.6.
1 follows from item 2 of Lemma˜4.7. 2 follows from the items 1 and 3 of Lemma˜4.7.
See 4.9
Proof A.7.
@@
  • Otherwise, let $z\in\mathcal{V}\mathsf{ar}(\pi)\cap(\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell))$. Then, we have $\gamma]\!]_{\hat{\delta}^{\prime}}(z)=[\![\gamma]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}(z)=[\![\gamma(z)]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}=[\![z]\!]_{\hat{\delta}[\vec{w}\mapsto\vec{v}]}=\{\hskip 0.85358pt\vec{w}\mapsto\vec{v}\hskip 0.85358pt\}(z)=\hat{\sigma}(z)$.
@@
See 5.4
Proof A.8.
[…]
See 5.5
Proof A.9.
The claim follows from (1) $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}\subseteq\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$, and (2) if $u,v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ with $u<v$, then $v\notin\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. (1) is immediate from Lemma˜5.3 3. To show (2), suppose $u,v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ with $u<v$, i.e., there exist substitutions $\sigma,\theta$ such that (3) $u=s\sigma$, (4) $\sigma(X)\subseteq\mathcal{V}\mathsf{al}$, (5) $\sigma\vDash\exists\vec{x}.\ \varphi$, (6) $v=u\theta$, and (7) $v\not\leqslant u$. Note that from (3)–(5) we obtain $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$, and thus $v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ follows by Lemma˜4.3 2 and 6. Now we suppose $v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ and derive a contradiction. Let $s=s[x_{1},\ldots,x_{m},y_{1},\ldots,y_{n}]_{p_{1},\ldots,p_{m},q_{1},\ldots,q_{n}}$ with $\mathcal{P}\mathsf{os}_{X}(s)=\{\hskip 0.85358ptp_{1},\ldots,p_{m}\hskip 0.85358pt\}$ and $\mathcal{P}\mathsf{os}_{\mathcal{V}\setminus X}(s)=\{\hskip 0.85358ptq_{1},\ldots,q_{n}\hskip 0.85358pt\}$. Then, by (3) and (4), $u=s[v_{1},\ldots,v_{m},y_{1}\sigma,\ldots,y_{n}\sigma]$ with $v_{1},\ldots,v_{m}\in\mathcal{V}\mathsf{al}$. By (6), $v=u\theta=s[v_{1},\ldots,v_{m},y_{1}\sigma\theta,\ldots,y_{n}\sigma\theta]$. From our assumption that $v\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$, it follows that the sequence of terms $\langle y_{1}\sigma\theta,\ldots,y_{n}\sigma\theta\rangle$ is a sequence of variables which is a renaming of $\langle y_{1},\ldots,y_{n}\rangle$. Hence, $v\leqslant u$. This is a contradiction to (7).
See 6.1
Proof A.10.
@@
    The specific substitution $\hat{\delta}$ in the proof of Lemma˜4.7 satisfies $\hat{\delta}(x)=x$ for $x\in\mathcal{V}\mathsf{ar}(s)\setminus X$, as $(\mathcal{V}\mathsf{ar}(\pi\gamma)\cup\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\setminus X)\cap\mathcal{V}\mathsf{ar}(s)=\varnothing$. Thus, it follows that $u=s\hat{\delta}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$.